Global PPF and Inereasing Oppertunity Costs The countries of England and Ihaly have the production possibilities tables for sheep and hogs s the costs of prodocing sheep and hogs Englands PRODUC TION POSSIBILITIES TABLE are constant in both countries. Assume there are only two countries in the world. itarys PROOUCTION POSSIBILITIES TABLE Aernelv 25 20 15 t0 Sheep Hogs 20 16 12 840 Hogs Taxk 8: If both England and Italy were to produce only sheep, how many sheep would be produced? Task 9: If both England and Italy were to produce only hogs, how many hogs would be produced? Task 10: If both England and Italy were to specialize and produce based on their comparative advantage, and if they are producing at full capacity, how many sheep and hogs will be produced?
Answer #1

Task 8.

England will produce 25 sheep and Italy will produce 20 sheep, totaling 45 sheep.

Task 9

England will produce 25 hogs and Italy will produce 15 hogs, totaling 40 hogs.

Task 10

Opportunity cost is the value of the next best alternative.

Sheep in terms of opportunity cost of hogs

Hogs in terms of opportunity cost of sheep.

England

25/25=1

25/25=1

Italy

15/20=0.75

20/15=1.33

Italy has a comparative advantage in the production of sheep since its opportunity is lower, 0.75 versus 1 for England.

England has a comparative advantage in the production of hogs since its opportunity is lower, 1 versus 1.33 for Italy.

Italy will specialize in the production of sheep and it will produce 20 sheep and England will specialize in the production of hogs and it will produce 25 hogs.

Engar PPF 205 タ ST l bs. 6只 0 1bs